re arrange the equation to make x the subject
ax +b =px +q
g (x +1)= f -x

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

re arrange the equation to make x the subject
ax +b =px +q
ax-px = q-b
x(a-p) = q-b
x = (q-b)/(a-p)
-----------------------
g (x +1)= f -x
qx + q = f - x
qx + x = f-q
(q+1)x = f-q
x = (f-q)/(q+1)
---------------------
Cheers,
Stan H.
------------

Answer by ewatrrr(24785)

ax +b =px +q
ax - px = q - b
x(a-p) = (q-b)
x = (q-b)/(a-p)
............
g (x +1)= f -x
gx + g = f - x
gx + x = (f-g)
x(g+1) = (f-g)
x = (f-g)/(g+1)